OpenStudy (faman39):

Lydia received a parking ticket because she parked 10 feet from a fire hydrant. She had estimated her distance from the fire hydrant as being 16 feet. What was her percent error?

OpenStudy (anonymous):

she was really 10 feet she thought 16 feet she was 6 feet off her error then is 6feet/10feet error/real value .6=60%

Directrix (directrix):

% error = | your result - accepted value | x 100 % accepted value ABS[ (16 - 10)/ 10 ] x 100 % = 6/10 x 100% = 60% error